arricandy:

What step comes first to solve 18 ÷ 2 - 5 + 7 - 12?

luvvdani1516:

Think of PEMDAS to solve this. P- parentheses E- exponents M-multiplication D- division A- adding S- subtracting Or you can just input it all into a calculator and hit equal to get the answer.

Mathyg:

18/2=9 -5+7=2 9-2=7 7-12=-5 Apply RULE: BODMAS (bracket open divide multiply addition subtration) 18/2-5+7-12 9-5+7-12 9+2-12 7-12 -5

luvvdani1516:

@mathyg wrote:
18/2=9 -5+7=2 9-2=7 7-12=-5 Apply RULE: BODMAS (bracket open divide multiply addition subtration) 18/2-5+7-12 9-5+7-12 9+2-12 7-12 -5
Try not to give the exact answer next time. You want the person whose asking to actually be able to do the work themselves.

Mathyg:

Oh is that so..

luvvdani1516:

Yeah you're not supposed to do all of the work for them

MadamMystery:

Firstly divide, secondly multiply, next add and at last subtract. There is a rule which can help you B-Brackets O-Of D- Division M-Multiplication A-Addition S-Subtraction

Astrid1:

Actually you can do the work for them as long as they know exactly what you're talking about. Make sure things are clear and that they understand.

MadamMystery:

@luvvdani1516 wrote:
Think of PEMDAS to solve this. P- parentheses E- exponents M-multiplication D- division A- adding S- subtracting Or you can just input it all into a calculator and hit equal to get the answer.
I guess division comes before multiplication...

omegataurus:

18 / 2 is the first step

Mathyg:

@madammystery wrote:
@luvvdani1516 wrote:
Think of PEMDAS to solve this. P- parentheses E- exponents M-multiplication D- division A- adding S- subtracting Or you can just input it all into a calculator and hit equal to get the answer.
I guess division comes before multiplication...
Yes u are right